(1 point) Find the solution to initial value problem dạy dt2 dy 169 + 64y = 0, y(0) dt = 10, y(0) = 4 The solution is
Answer #1

initial value broblem i Solution Given dy 4+2 -15 day E4y = with -(1) yo)-10, y Co=4 which can be written as (D?-16D +64)450